A projectile is fired in such way that its horizontal range is equal to three times its maximum height. What is the angle projection?
1
Expert's answer
2021-08-09T06:51:21-0400

Given:

"R=3H"

The range

"R=\frac{v_0^2\sin2\theta}{g}"

the maximum height

"H=\frac{v_0^2\sin^2\theta}{2g}"

Hence,

"\frac{v_0^2\sin2\theta}{g}=3*\frac{v_0^2\sin^2\theta}{2g}"

"\sin2\theta=3/2\sin^2\theta"

"2\sin\theta\cos\theta=3/2\sin^2\theta"

"\tan\theta=4/3"

"\theta=53^{\circ}"